OpenStudy (anonymous):

how to solve log 5+log x=log 10

OpenStudy (anonymous):

Use: \[\log(a) + \log(b) = \log(ab)\]

OpenStudy (anonymous):

I guess this is the log to base ten, often written as lg, so isolate the x and then 10^

OpenStudy (anonymous):

so log 10 will each equal log(5x)

OpenStudy (anonymous):

So: \[\log(5x) = \log(10) \implies 5x = 10 \implies \color{green}{x = 2}\]

OpenStudy (anonymous):

gracias ustedes

OpenStudy (anonymous):

Now you can cancel log on both the sides or you can say taking Anti Log both the sides we get: 5x = 10 Now solve for x here..

OpenStudy (anonymous):

thank you i understand now

OpenStudy (anonymous):

gracias ustedes What does that mean?? Thank you???

OpenStudy (anonymous):

it means thank y'all

OpenStudy (anonymous):

Oh I see.. Welcome dear..

OpenStudy (anonymous):

By the way which language is this??